OpenStudy (anonymous):

find the principal roots indicated: \[\sqrt{121x^6y^4}\]

OpenStudy (anonymous):

\[\sqrt{121x^6y^4}\]

OpenStudy (mertsj):

Do you know that 121 = 11 x 11? And x^6=(x^3)(x^3) And y^4=(y^2)(y^2) ?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Well then, what is: \[\sqrt{121}\]

OpenStudy (anonymous):

11

OpenStudy (mertsj):

What is \[\sqrt{x^6}\]

OpenStudy (anonymous):

oh,,, x^3 why is it always so easy when you say it???

OpenStudy (mertsj):

Don't know.